The following data were drawn from the records of Solomon Corporation. Planned volume for year (static budget) Standard direcComplete this question by entering your answers in the tabs below. Req A Red B ReqC Req D Req E to G Prepare a labor varianceComplete this question by entering your answers in the tabs below. Req A Req B Reqc Req D Req E to G Calculate the labor pricComplete this question by entering your answers in the tabs below. Req A Req B Reqc Req D Req E to G muone verance, indicateComplete this question by entering your answers in the tabs below. Req A Req B Reqc Req D Req E to G Calculate the materials
Answer #1

As per HOMEWORKLIB RULES, we can answer only 4 sub questions in a single question, i am answering the first 4 questions, this is for quality ensurance purpose, student may re upload the remaining questions.

  1. Material varience information table and variences,
    1. Material varience information table qty/amt
      Standard price 1.6 Per pound
      Actual price 2.0 Per pound
      Standard quantity for flexible budget 18,000* Pounds
      Actual quantity used 16,200** Pounds

      * 4,500 units * 4 pounds per unit, ** 4,500 units * 3.6 pounds per unit

    2. Varience F/U
      Material price varience* 6,480 U
      Material usage varience** 3,600 F

      *  (Standard unit price - Actual unit price) * Actual Quantity used = (2-1.6)*16,200. **  (Standard usage - Actual usage) x Standard price per unit = (18000 - 16200) * 2.

  2. Labour varience information table and labour variences,

    1. Labour varience information table qty/amt
      Standard price 5.0 Per pound
      Actual price 4.6 Per pound
      Standard hours for flexible budget* 9,450 Pounds
      Actual hours** 10,350 Pounds

      * 4500 units * 2.1 hours = 9,450 ** 4,500 units * 2.3 = 10,350

    2. Varience F/U
      Labour price varience* 4,140 F
      Labour usage varience** 4,500 U

      * (Standard price - actual price)*actual hours = (5-4.6)*10,350 ** (Standard hours - actual hours)*standard price = (9,450 - 10,350) * 5